The Bogart Company produces 6,000 units of item SLM 46 annually at a total cost of $240,000.

Direct materials $ 24,000
Direct labor 66,000
Variable overhead 54,000
Fixed overhead 96,000
Total $ 240,000


The Conner Company has offered to supply all 6,000 units of SLM 46 per year for $35 per unit. If Bogart accepts the offer, $8 per unit of the fixed overhead would be saved. In addition, some of Bogart's leased facilities could be vacated, reducing lease payments by $96,000 per year. At what price would Bogart be indifferent to Conner's offer?
Answer #1

Relevant cost = direct material + direct labor + variable overhead + Avoidable fixed cost + lease rent

= 24000+66000+54000+48000 + 96000

= $288000

Indifferent price = 288000/6000 = $48 per unit